# DecoR: Deconfounding Time Series with Robust Regression

### v1

@@ 4.1 Example: Band-Limited Processes
and analogously for $U$. Let $c\coloneqq|S_{X}\cap S_{U}|$ and assume that $|S_{X}|\geq 2c+1$. Assume further that, there exists $r>0$ such that for all $k\in S_{X}$ it holds that almost surely $|a_{k}^{U}|\geq r$. This implies Assumption 2 (ii) and therefore, by Theorem 4.1, DecoR-BFS is consistent. If, additionally, condition (10) is satisfied – this is the case with high probability, for example, if $\{a_{k}^{X}\}_{k\in S_{X}}$ are i.i.d. Gaussian and $|S_{X}|$ is large – then, by Theorem 4.2, DecoR-Tor is also consistent. Moreover, OLS can be inconsistent in this setting, as shown in the following proposition.
Proposition 4.3.
Assume Setting 1 and assume Assumption 2 holds. Then, $\hat{\beta}^{\phi,n}_{\texttt{OLS}}$ is not necessarily a consistent estimator for $\beta$. Even if $X$ is a band-limited process, OLS may be inconsistent.
We regard these results surprising in that they describe a setting, in which the number of outliers increases linearly in $n$ and there is a constant [^9] The noise variance is only constant in the time domain. In the frequency domain, it vanishes with increasing $n$. noise variance, and still robust regression yields a consistent estimator (and OLS does not). As discussed in Appendix E.3 such an estimator does not exist for i.i.d. data.
@@ 5 Empirical Evaluation
We validate our theoretical findings through experiments on synthetic data. In the experiments, we generate data according to the model described in Setting 1 and set $X\coloneqq U+\epsilon_{X}$. We sample $\epsilon_{X}$ and $U$ from either two independent Ornstein-Uhlenbeck processes or two independent band-limited processes. For the orthogonal basis $\phi$, we consider both the cosine basis (see Definition C.1) and the Haar basis (see Definition C.2). The noise in $Y$ is set to have variance $\sigma_{\eta}^{2}\in\{0,1,4\}$. To assess the accuracy of the different approaches, we calculate the mean absolute prediction error: $\mathrm{MAE}\coloneqq\sum_{i=1}^{n}|\beta_{i}-\hat{\beta}^{n}_{i}|/n$. For the remainder of this section, if not specified otherwise, we assume independent band-limited processes, the cosine basis, noise variance of $\sigma^{2}_{\eta}=1$, fraction of outliers of $25\%$ and $a=0.7$. Detailed information on the experimental setup is available in Appendix F.2 and code has been made available at https://github.com/fschur/robust_deconfounding.
@@
We first compare the average estimation error of OLS, DecoR-Tor and DecoR-BFS. As discussed in Section 3, the computation time of DecoR-BFS scales exponentially with the sample size $n$, so we only consider sample sizes up to $n=16$. Table 1 shows the results. In the absence of noise, that is, if $\sigma^{2}_{\eta}=0$, DecoR-BFS correctly point-identifies the true parameter $\beta$ for all samples sizes – this is in agreement with Theorem 4.1. Although the average estimation error of DecoR-Tor decreases as the sample size increases, it does not reach zero, unlike the estimation error of DecoR-BFS. This discrepancy arises because condition (10) is often not satisfied in smaller sample sizes (not shown). The discrepancy also suggests a trade-off between computational costs and average estimation error. Since DecoR is modular in that it can be applied with any robust regression technique, it directly benefits from any methodological advancement in the field of robust regression. As expected, the estimation error for OLS remains constant regardless of sample size and is significantly larger than that of both DecoR-Tor and DecoR-BFS.
In Section 4.1 we have shown that under suitable conditions DecoR is consistent when $U$ and $X$ are band-limited processes. Figure 3 (left) shows that the mean absolute estimation error for processes from this class vanishes with growing sample size, supporting considers Ornstein-Uhlenbeck processes and suggests that DecoR might also be consistent for this class of processes. We show, however, in Appendix E.3 that there are settings with Ornstein-Uhlenbeck processes where no consistent estimator exists; it therefore seems that many of the cases in Figure 3 (right) contain confounding that is not worst-case, so DecoR is still consistent. In both cases (Figure 3 (left) and Figure 3 (right)) the error of OLS does not seem to vanish with growing sample size.
@@ 6 Summary and Future Work
Looking ahead, we see several promising directions for future work. First, adapting our results to datasets where the data points are not observed in regular intervals. Similarly, extending our framework to include the asymptotics of longer observational time horizons, as opposed to shorter time intervals, would be interesting. Second, exploring scenarios where the effect of the hidden confounder has a linear and sparse effect on $X$, but is dense toward $Y$, may be relevant for practical applications. For such cases, similarly as before, we have
$T_{k}^{\phi,n}(Y)=T_{k}^{\phi,n}(X)^{\top}\beta+T_{k}^{\phi,n}(U)+T_{k}^{\phi,n}(\eta).$
However, in contrast to Setting 1, the hidden confounder does not disappear, that is, $(T_{k}^{\phi,n}(U))_{G^{c}}\neq 0$. Instead, assuming $X=\alpha U+\epsilon_{X}$ for some $\alpha\in\mathbb{R}^{d}$ and $\epsilon_{X}\perp\!\!\!\perp X$, one may be able to exploit that $(T_{k}^{\phi,n}(X))_{G_{n}^{c}}^{\top}(T_{k}^{\phi,n}(U))_{G_{n}^{c}}\to 0$. Lastly, it could be interesting to extend the results about the lower bound to other noise distributions such as the Gaussian distribution. We hypothesize that, in this case, our approach would be consistent under minor modifications. Third, another promising direction is to consider nonlinear causal effects. In this setting, one strategy could be to generalize the derived results to finite-dimensional feature transformations or to kernelize the approach.
@@ Acknowledgements
We thank Nicolai Meinshausen for his helpful input regarding the idea of sparse confounding in the frequency domain and Michael Law for his helpful comments on convergence of the OLS estimator for time series data. We have used ChatGPT to improve minor language formulations.
@@ Appendix C Additional Definitions
Definition C.1 (Cosine basis).
@@
Definition C.2 (Haar basis (Haar, 1909)).
@@
Definition C.3 (Bachmann–Landau notation).
@@
Definition C.4.
Let $X_{1},X_{2},\ldots$ be a sequence of random variables and $g:\mathbb{N}\to\mathbb{R}$ a real-valued function. We say that with high probability it holds that
$X_{n}\in\mathcal{O}(g(n))$
@@
We define $X_{n}\in o(g(n))$ with high probability and $X_{n}\in\Omega(g(n))$ with high probability analogously.
@@ D.1 Lemmas
Lemma D.1 (Theorem 6.3.2 of Vershynin (2018)).
@@
We are now able to prove Theorem D.2. Let
@@
This implies by Lemma D.3 that
@@
Proof of Theorem 3.1.
Let $U_{n}\in\mathrm{Inl}(\mathcal{U}_{n})$. By the proof of Theorem D.2, specifically (15) and (16), we have that
@@
By Lemma D.1 there exists $K>0$ such that for all $t\geq 0$ we have
@@
As shown in the proof of Theorem D.2, we have
@@
D.3 Proof of Lemma 3.2
The number of distinct sets $S$ that Torrent can output is trivially bounded by $\binom{n}{\lfloor a\rfloor}$. Furthermore, it holds that
$\left\lVert X_{S_{t+1}}(\beta-\hat{\beta}^{t+1}_{\texttt{Tor}})+\epsilon_{S_{t+1}}+o_{S_{t+1}}\right\rVert_{2}$ $\leq\left\lVert X_{S_{t+1}}(\beta-\hat{\beta}^{t}_{\texttt{Tor}})+\epsilon_{S_{t+1}}+o_{S_{t+1}}\right\rVert_{2}$
$\leq\left\lVert X_{S_{t}}(\beta-\hat{\beta}^{t}_{\texttt{Tor}})+\epsilon_{S_{t}}+o_{S_{t}}\right\rVert_{2}.$
The first inequality holds because $\hat{\beta}_{\texttt{Tor}}^{t+1}$ is the least squares estimate for data $X_{S_{t+1}}$, while the second holds because of the hard thresholding step. Furthermore, this implies that Torrent converges after finitely many steps.
D.4 Proof of Theorem 3.3
Assume that $X_{S_{t}}^{\top}X_{S_{t}}$ is invertible (otherwise, the bound is void). We begin by observing that the optimality of the model $\hat{\beta}^{t}_{\texttt{Tor}}$ on the active set $S_{t}$ ensures
$\begin{split}\left\lVert\hat{\beta}^{t}_{\texttt{Tor}}-\beta\right\rVert_{2}&=\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}\left(X_{S_{t}}^{\top}\epsilon_{S_{t}}+X_{S_{t}}^{\top}o_{S_{t}}\right)\right\rVert_{2}\\ &\leq\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\right\rVert_{2}\left\lVert o_{S_{t}}\right\rVert_{2}.\end{split}$ (20)
@@
$\begin{split}\left\|X_{S_{t+1}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)+\epsilon_{S_{t+1}}+o_{S_{t+1}}\right\|_{2}^{2}&=\left\|Y_{S_{t+1}}-X_{S_{t+1}}\hat{\beta}^{t}_{\texttt{Tor}}\right\|_{2}^{2}\\ &\leq\left\|Y_{G_{n}^{\mathsf{c}}}-X_{G_{n}^{\mathsf{c}}}\hat{\beta}^{t}_{\texttt{Tor}}\right\|^{2}_{2}\\ &=\left\|X_{G_{n}^{\mathsf{c}}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)+\epsilon_{G_{n}^{\mathsf{c}}}\right\|_{2}^{2}.\end{split}$ (21)
Let $H_{t+1}\coloneqq S_{t+1}\backslash G_{n}^{\mathsf{c}}$ and $M_{t+1}\coloneqq G_{n}^{\mathsf{c}}\backslash S_{t+1}$. It then follows from (21)
$\left\|X_{H_{t+1}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)+\epsilon_{H_{t+1}}+o_{H_{t+1}}\right\|_{2}\leq\left\|X_{M_{t+1}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)+\epsilon_{M_{t+1}}\right\|_{2}.$
@@
Let $V_{t+1}\coloneqq H_{t+1}\cup M_{t+1}$ then $V_{t+1}=(S_{t+1}\cup G_{n}^{\mathsf{c}})\setminus(G_{n}^{\mathsf{c}}\cap S_{t+1})=V(S_{t})$. An application of the triangle inequality and the fact that $\left\|o_{H_{t+1}}\right\|_{2}=\left\|o_{S_{t+1}}\right\|_{2}$ gives us
@@
$\begin{split}\left\|o_{S_{t+1}}\right\|_{2}&\leq\left\|X_{M_{t+1}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)\right\|_{2}+\left\|X_{H_{t+1}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)\right\|_{2}+\left\|\epsilon_{H_{t+1}}\right\|_{2}+\left\|\epsilon_{M_{t+1}}\right\|_{2}\\ &\leq\sqrt{2}\left\lVert X_{V_{t+1}}\right\rVert_{2}\left\|\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right\|_{2}+\sqrt{2}\|\epsilon_{V_{t+1}}\|_{2}\\ &\leq\sqrt{2}\left\lVert X_{V_{t+1}}\right\rVert_{2}\left(\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\right\rVert_{2}\left\lVert o_{S_{t}}\right\rVert_{2}\right)\\ &\quad+\sqrt{2}\|\epsilon_{V_{t+1}}\|_{2}\\ &=\sqrt{2}\left\lVert X_{V_{t+1}}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\right\rVert_{2}\left\lVert o_{S_{t}}\right\rVert_{2}+\sqrt{2}\left\lVert X_{V_{t+1}}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}\\ &\quad+\sqrt{2}\|\epsilon_{V_{t+1}}\|_{2}.\end{split}$ (22)
@@
Here the second inequality follows since for all $x\in\mathbb{R}^{d}$ we have $\left\lVert x\right\rVert_{1}\leq\sqrt{d}\left\lVert x\right\rVert_{2}$. Let $t$ be the iteration where Torrent has converged (see Lemma 3.2). We prove that we can assume that either $S_{t}=S_{t+1}$ or $S_{t+1}=S_{t+2}$ hold. First, assume that $\hat{\beta}^{t+1}_{\texttt{Tor}}=\hat{\beta}^{t}_{\texttt{Tor}}$. Then, by Algorithm 3 we have $S_{t+1}=S_{t+2}$. Next, assume $\hat{\beta}^{t+1}_{\texttt{Tor}}\neq\hat{\beta}^{t}_{\texttt{Tor}}$. If $X_{S_{t+1}}^{T}X_{S_{t+1}}$ is not invertible, the bound is void and there is nothing to show. If $X_{S_{t+1}}^{T}X_{S_{t+1}}$ is invertible, then $\hat{\beta}^{t+1}_{\texttt{Tor}}$ is the unique minima and the algorithm did not stop, which is a contraction to the assumption that $S_{t}$ is the output of Torrent. Without loss of generality, we can therefore assume that $S_{t}=S_{t+1}$.
@@
Together with $S_{t}=S_{t+1}$ this yields
$\left\|o_{S_{t}}\right\|_{2}\leq\sqrt{2}\eta\left\lVert o_{S_{t}}\right\rVert_{2}+2\left\lVert X_{V_{t}}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\sqrt{2}\|\epsilon_{V_{t}}\|_{2}$
@@
$\left\|o_{S_{t}}\right\|_{2}\leq\frac{2\left\lVert X_{V_{t}}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\sqrt{2}\|\epsilon_{V_{t}}\|_{2}}{1-\sqrt{2}\eta}.$
@@
$\quad+\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\right\rVert_{2}\frac{2\left\lVert X_{V_{t}}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\sqrt{2}\|\epsilon_{V_{t}}\|_{2}}{1-\sqrt{2}\eta}.$
D.5 Proof of Corollary 3.4

### v2

@@ 4.1 Example: Band-Limited Processes
and analogously for $U$. Let $c\coloneqq|S_{X}\cap S_{U}|$ and assume that $|S_{X}|\geq 2c+1$. Assume further that, there exists $r>0$ such that for all $k\in S_{X}$ it holds that almost surely $|a_{k}^{U}|\geq r$. This implies Assumption 2 (ii) and therefore, by Theorem 3, DecoR-BFS is consistent. If, additionally, condition (11) is satisfied – this is the case with high probability, for example, if $\{a_{k}^{X}\}_{k\in S_{X}}$ are i.i.d. Gaussian and $|S_{X}|$ is large – then, by Theorem 4, DecoR-Tor is also consistent. Moreover, OLS can be inconsistent in this setting, as shown in the following proposition.
Proposition 1.
Assume Setting 1 and assume Assumption 2 holds. Then $\hat{\beta}^{\phi,n}_{\texttt{OLS}}$ is not necessarily a consistent estimator for $\beta$. Even if $X$ is a band-limited process, OLS may be inconsistent.
We regard these results surprising in that they describe a setting in which the number of outliers increases linearly in $n$ and there is a constant [^9] The noise variance is only constant in the time domain. In the frequency domain, it vanishes with increasing $n$. noise variance, and still robust regression yields a consistent estimator (and OLS does not). As discussed in Appendix E.3 such an estimator does not exist for i.i.d. data.
@@ 5 Empirical Evaluation
We validate our theoretical findings through experiments on synthetic data and real-world data. The data sets and code for these experiments have been made available at https://github.com/fschur/robust_deconfounding.
5.1 Synthetic Data
In our synthetic experiments, we generate data according to the model described in Setting 1 and set $X\coloneqq U+\epsilon_{X}$. We sample $\epsilon_{X}$ and $U$ from either two independent Ornstein-Uhlenbeck processes or two independent band-limited processes. For the orthogonal basis $\phi$, we consider both the cosine basis (see Definition 2) and the Haar basis (see Definition 3). The noise in $Y$ is set to have variance $\sigma_{\eta}^{2}\in\{0,1,4\}$. To assess the accuracy of the different approaches, we calculate the mean absolute prediction error: $\mathrm{MAE}\coloneqq\sum_{i=1}^{n}|\beta_{i}-\hat{\beta}^{n}_{i}|/n$. For the remainder of this section, if not specified otherwise, we assume independent band-limited processes, the cosine basis, noise variance of $\sigma^{2}_{\eta}=1$, fraction of outliers of $25\%$ and $a=0.7$. Detailed information on the experimental setup is available in Appendix F.1.
@@ 5.1 Synthetic Data
We first compare the average estimation error of OLS, DecoR-Tor and DecoR-BFS. As discussed in Section 3, the computation time of DecoR-BFS scales exponentially with the sample size $n$, so we only consider sample sizes up to $n=16$. Table 1 shows the results. In the absence of noise, that is, if $\sigma^{2}_{\eta}=0$, DecoR-BFS correctly point-identifies the true parameter $\beta$ for all samples sizes – this is in agreement with Theorem 3. Although the average estimation error of DecoR-Tor decreases as the sample size increases, it does not reach zero, unlike the estimation error of DecoR-BFS. This discrepancy arises because condition (11) is often not satisfied in smaller sample sizes (not shown). The discrepancy also suggests a trade-off between computational costs and average estimation error. Since DecoR is modular in that it can be applied with any robust regression technique, it directly benefits from any methodological advancement in the field of robust regression. As expected, the estimation error for OLS remains constant regardless of sample size and is significantly larger than that of both DecoR-Tor and DecoR-BFS.
In Section 4.1 we have shown that under suitable conditions DecoR is consistent when $U$ and $X$ are band-limited processes. Figure 3 (left) shows that the mean absolute estimation error for processes from this class vanishes with increasing sample size, supporting this result. Figure 3 (right) considers Ornstein-Uhlenbeck processes and suggests that DecoR might also be consistent for this class of processes. The results of Appendix E.3, however, suggest that there are settings with Ornstein-Uhlenbeck processes where no consistent estimator exists; it therefore seems that many of the cases in Figure 3 (right) contain confounding that is not worst-case, so DecoR is still consistent. In both cases (Figure 3 (left) and Figure 3 (right)) the error of OLS does not seem to vanish with growing sample size.
@@
5.2 Application to Earth System Science
An important challenge in Earth System Science is to distinguish externally forced changes in climate systems from internal variability or “climate noise” (Madden, 1976; Schneider and Kinter, 1994). Drawing inspiration from Sippel et al. (2019) we consider a specific instance involving atmospheric circulation and precipitation and frame the problem as a hidden confounder problem. Specifically, in our study, the covariate time series $X$ represents atmospheric circulation, in this case pressure at sea level, the target time series $Y$ corresponds to precipitation, and the hidden confounder $U$ denotes external forcing (such as greenhouse gas concentrations), which we presume to be sparse in the frequency domain. The goal is to estimate the internal circulation-induced contribution to variability of precipitation. Thus, after estimating the causal effect of $X$ on $Y$ we remove the confounded frequencies from $X$ and use the causal function from $X$ to $Y$ to produce fitted values; these fitted values correspond to the internal circulation-induced contribution to the variability of precipitation (assuming that the confounder acts linearly on $X$); the corresponding residuals of $Y$ can be used as an estimate of the total forced response. We use the data set provided by the ClimEx project (Leduc et al., 2019), based on a single run of a climate model, and consider precipitation data during the winter months (December, January, and February) in the Alps and pressure at sea level data during the same time over Europe. Our method, DecoR-Tor with parameter $a=0.9n$ and a cosine basis, is compared to the approach developed by Sippel et al. (2019). A critical aspect of the approach of Sippel et al. (2019) is the prerequisite that the frequency bands in which the confounder is active are known; more specifically, one may assume that the confounder (external forcing) is smooth (see also Wallace et al. (1995)), that is, the external forcing consists solely of low-frequency components. If this assumption approximately holds, we expect DecoR to exclude mainly low-frequency components when applied to the data set.
Figure 4: Histogram of the frequencies excluded by DecoR in the final iteration (left) and centred average 5-year precipitation in the alps over time (right); here, the methods’ outputs estimate the precipitation without external forcing. The “ground truth” data is the precipitation provided by the ClimEx project (Leduc et al., 2019).
Indeed, an analysis of the frequencies excluded by DecoR, shown in Figure 4(a), confirms that our algorithm predominantly filters out the confounded low frequencies rather than the unconfounded high frequencies. Furthermore, the coefficient of determination ($R^{2}$) for the fitted values (see above) DecoR equals $0.76$, the same as the one obtained by the method of Sippel et al. (2019). Thus, DecoR seems competitive even without using the expert knowledge of which frequencies to exclude. The fitted values of the precipitation patterns (which ignore the confounded frequencies) are depicted in Figure 4(b); both DecoR and the method used by Sippel et al. (2019) estimate lower mean precipitation than observed in the data set, (which is to be expected, as it is believed that external forcing induces an upward trend in the data (Foster and Rahmstorf, 2011; Deser et al., 2016)). In conclusion, our method matches the effectiveness of the approach by Sippel et al. (2019) while imposing fewer assumptions.
@@ 6 Summary and Future Work
Looking ahead, we see several promising directions for future work. First, adapting our results to data sets where the data points are not observed in regular intervals. Similarly, extending our framework to include the asymptotics of longer observational time horizons, as opposed to shorter time intervals, would be interesting. Second, exploring scenarios where the effect of the hidden confounder has a sparse effect on $X$, but is dense toward $Y$, may be relevant for practical applications. We hypothesize that, in this case, our approach would be consistent under minor modifications. Third, it could be interesting to extend the results about the lower bound to other noise distributions such as the Gaussian distribution. Fourth, another promising direction is to consider nonlinear causal effects. In this setting, one strategy could be to generalize the derived results to finite-dimensional feature transformations or to kernelize the approach.
@@ Acknowledgements
We thank Nicolai Meinshausen for his helpful input regarding the ideas of sparse confounding in the frequency domain and the real data application and Michael Law for his helpful comments on convergence of the OLS estimator for time series data. We thank Maybritt Schillinger for helping us with the Earth System Science experiment. We also thank Pio Blieske for taking the time to read this paper and offering invaluable feedback. We have used ChatGPT to improve minor language formulations.
@@ Appendix C Additional Definitions
Definition 2 (Cosine basis).
@@
Definition 3 (Haar basis (Haar, 1909)).
@@
Definition 4 (Bachmann–Landau notation).
@@
Definition 5.
Let $X_{1},X_{2},\ldots$ be a sequence of random variables and $g:\mathbb{N}\to\mathbb{R}$ a real-valued function. We say that we have
$X_{n}\in\mathcal{O}_{\mathbb{P}}(g(n))$
@@
We define $X_{n}\in o_{\mathbb{P}}(g(n))$ and $X_{n}\in\Omega_{\mathbb{P}}(g(n))$ analogously.
@@ D.1 Lemmas
Lemma 2 (Theorem 6.3.2 of Vershynin (2018)).
@@ D.2 Proof of Theorem 1
We are now able to prove Theorem 5. Let
@@
This implies by Lemma 3 that
@@
Proof of Theorem 1.
Let $U_{n}\in\mathrm{Inl}(\mathcal{U}_{n})$. By the proof of Theorem 5, specifically (16) and (17), we have that
@@
By Lemma 2 there exists $K>0$ such that for all $t\geq 0$ we have
@@
As shown in the proof of Theorem 5, we have
@@
D.3 Proof of Lemma 1
The number of distinct sets $S$ that Torrent can output is trivially bounded by $\binom{n}{\lfloor a\rfloor}$. Furthermore, it holds for all $t$ that
$\left\lVert X_{S_{t+1}}(\beta-\hat{\beta}^{t+1}_{\texttt{Tor}})+\epsilon_{S_{t+1}}+o_{S_{t+1}}\right\rVert_{2}$ $\leq\left\lVert X_{S_{t}}(\beta-\hat{\beta}^{t+1}_{\texttt{Tor}})+\epsilon_{S_{t}}+o_{S_{t}}\right\rVert_{2}$ (21)
$\leq\left\lVert X_{S_{t}}(\beta-\hat{\beta}^{t}_{\texttt{Tor}})+\epsilon_{S_{t}}+o_{S_{t}}\right\rVert_{2}.$ (22)
The first inequality holds because of the hard thresholding step, while the second holds because $\hat{\beta}_{\texttt{Tor}}^{t+1}$ is the least squares estimate for data $X_{S_{t}}$. This implies that Torrent converges after finitely many steps.
D.4 Proof of Theorem 2
Assume that $X_{S_{t-1}}^{\top}X_{S_{t-1}}$ is invertible (otherwise, the assumptions for $\eta$ are not satisfied). We begin by observing that the optimality of the model $\hat{\beta}^{t}_{\texttt{Tor}}$ on the active set $S_{t}$ ensures
$\begin{split}\left\lVert\hat{\beta}^{t}_{\texttt{Tor}}-\beta\right\rVert_{2}&=\left\lVert\left(X_{S_{t-1}}^{\top}X_{S_{t-1}}\right)^{-1}\left(X_{S_{t-1}}^{\top}\epsilon_{S_{t-1}}+X_{S_{t-1}}^{\top}o_{S_{t-1}}\right)\right\rVert_{2}\\ &\leq\left\lVert\left(X_{S_{t-1}}^{\top}X_{S_{t-1}}\right)^{-1}X_{S_{t-1}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\left\lVert\left(X_{S_{t-1}}^{\top}X_{S_{t-1}}\right)^{-1}X_{S_{t-1}}^{\top}\right\rVert_{2}\left\lVert o_{S_{t-1}}\right\rVert_{2}.\end{split}$ (23)
@@ D.4 Proof of Theorem 2
$\begin{split}\left\|X_{S_{t}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)+\epsilon_{S_{t}}+o_{S_{t}}\right\|_{2}^{2}&=\left\|Y_{S_{t}}-X_{S_{t}}\hat{\beta}^{t}_{\texttt{Tor}}\right\|_{2}^{2}\\ &\leq\left\|Y_{G_{n}^{\mathsf{c}}}-X_{G_{n}^{\mathsf{c}}}\hat{\beta}^{t}_{\texttt{Tor}}\right\|^{2}_{2}\\ &=\left\|X_{G_{n}^{\mathsf{c}}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)+\epsilon_{G_{n}^{\mathsf{c}}}\right\|_{2}^{2}.\end{split}$ (24)
Let $H_{t}\coloneqq S_{t}\backslash G_{n}^{\mathsf{c}}$ and $M_{t}\coloneqq G_{n}^{\mathsf{c}}\backslash S_{t}$. It then follows from (24)
$\left\|X_{H_{t}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)+\epsilon_{H_{t}}+o_{H_{t}}\right\|_{2}\leq\left\|X_{M_{t}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)+\epsilon_{M_{t}}\right\|_{2}.$
@@
Let $V_{t}\coloneqq H_{t}\cup M_{t}$ then $V_{t}=(S_{t}\cup G_{n}^{\mathsf{c}})\setminus(G_{n}^{\mathsf{c}}\cap S_{t})=V(S_{t})$. An application of the triangle inequality and the fact that $\left\|o_{H_{t}}\right\|_{2}=\left\|o_{S_{t}}\right\|_{2}$ gives us
@@
$\begin{split}\left\|o_{S_{t}}\right\|_{2}&\leq\left\|X_{M_{t}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)\right\|_{2}+\left\|X_{H_{t}}\left(\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right)\right\|_{2}+\left\|\epsilon_{H_{t}}\right\|_{2}+\left\|\epsilon_{M_{t}}\right\|_{2}\\ &\leq\sqrt{2}\left\lVert X_{V_{t}}\right\rVert_{2}\left\|\beta-\hat{\beta}^{t}_{\texttt{Tor}}\right\|_{2}+\sqrt{2}\|\epsilon_{V_{t}}\|_{2}\\ &\leq\sqrt{2}\left\lVert X_{V_{t}}\right\rVert_{2}\left(\left\lVert\left(X_{S_{t-1}}^{\top}X_{S_{t-1}}\right)^{-1}X_{S_{t-1}}^{\top}\epsilon_{S_{t-1}}\right\rVert_{2}+\left\lVert\left(X_{S_{t-1}}^{\top}X_{S_{t-1}}\right)^{-1}X_{S_{t-1}}^{\top}\right\rVert_{2}\left\lVert o_{S_{t-1}}\right\rVert_{2}\right)\\ &\quad+\sqrt{2}\|\epsilon_{V_{t}}\|_{2}\\ &=\sqrt{2}\left\lVert X_{V_{t}}\right\rVert_{2}\left\lVert\left(X_{S_{t-1}}^{\top}X_{S_{t-1}}\right)^{-1}X_{S_{t-1}}^{\top}\right\rVert_{2}\left\lVert o_{S_{t-1}}\right\rVert_{2}+\sqrt{2}\left\lVert X_{V_{t}}\right\rVert_{2}\left\lVert\left(X_{S_{t-1}}^{\top}X_{S_{t-1}}\right)^{-1}X_{S_{t-1}}^{\top}\epsilon_{S_{t-1}}\right\rVert_{2}\\ &\quad+\sqrt{2}\|\epsilon_{V_{t}}\|_{2}.\end{split}$ (25)
@@
Here the second inequality follows since for all $a,b\in\mathbb{R}$ we have $(|a|+|b|)^{2}\leq 2(a^{2}+b^{2})$. Let $t$ be the iteration where Torrent has converged (see Lemma 1). We prove that we can assume that $S_{t}=S_{t-1}$. First, assume that $\hat{\beta}^{t}_{\texttt{Tor}}=\hat{\beta}^{t-1}_{\texttt{Tor}}$. Then, by the (deterministic) hard thresholding step of Algorithm 3 we have $S_{t}=S_{t-1}$. Next, assume $\hat{\beta}^{t}_{\texttt{Tor}}\neq\hat{\beta}^{t-1}_{\texttt{Tor}}$. If $X_{S_{t-1}}^{T}X_{S_{t-1}}$ is not invertible, the assumptions on $\eta$ are not satisfied. If $X_{S_{t-1}}^{T}X_{S_{t-1}}$ is invertible, then $\hat{\beta}^{t}_{\texttt{Tor}}$ is the unique minimum and, therefore, (21) is a strict inequality. Together with (22) this implies that the algorithm did not stop, which is a contraction to the assumption that $S_{t}$ is the output of Torrent. Without loss of generality, we can therefore assume that $S_{t}=S_{t-1}$.
@@
Together with $S_{t}=S_{t-1}$ and (25) this yields
$\left\|o_{S_{t}}\right\|_{2}\leq\sqrt{2}\eta\left\lVert o_{S_{t}}\right\rVert_{2}+\sqrt{2}\left\lVert X_{V_{t}}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\sqrt{2}\|\epsilon_{V_{t}}\|_{2}$
@@
$\left\|o_{S_{t}}\right\|_{2}\leq\frac{\sqrt{2}\left\lVert X_{V_{t}}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\sqrt{2}\|\epsilon_{V_{t}}\|_{2}}{1-\sqrt{2}\eta}.$
@@
$\quad+\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\right\rVert_{2}\frac{\sqrt{2}\left\lVert X_{V_{t}}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\sqrt{2}\|\epsilon_{V_{t}}\|_{2}}{1-\sqrt{2}\eta}.$
D.5 Proof of Corollary 1